OpenStudy (anonymous):

How do I solve this system of linear equation using the substitution method? 1/2x- 2/3y= -1 3/7x + y= 18

OpenStudy (anonymous):

on the second equation, transpose first the X Y should remain on the left side

OpenStudy (anonymous):

then you have now the value for Y, try substituting the value of Y on your second equation to the first equation

OpenStudy (anonymous):

Any ideas on how to clear the fraction?

OpenStudy (anonymous):

multiply both side by the LCD the LCD is 6, right?

OpenStudy (anonymous):

that will give you 3x-4y=-6

OpenStudy (anonymous):

on the second equation, multiply both sides by 7, to remove the fraction

OpenStudy (anonymous):

the new equation is 3x-4y=-6 3x+7y=126

OpenStudy (anonymous):

Yup. So I reworked the problem with the tips that you gave. 3x -4(126-3)= -6 3x-504+12= -6 3x-492= -6 3x=486 x= 162 3(162) - 4y= -6 486- 4y= -6 -4y= -492 y=123 3(162)- 4(123)= -6 -6=-6

OpenStudy (anonymous):

alright so you've got it right based on my computations x=-182/11 y= -120/11

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

okay, no problem, you are welcome!

OpenStudy (anonymous):

x=14 y=12

OpenStudy (anonymous):

@hamidx19 did I work the problem corruptly?

OpenStudy (anonymous):

1/2x- 2/3y= -1 3/7x + y= 18 equation 1 divided by -2/3: y=3/2+3/4x subbing into eq2: 3/7x+3/4x+3/2=18 33/28x=16.5 x=16.5/(33/28)=14 subbing back into eq1: 7-2/3y=-1, 2/3y=8, y=12

OpenStudy (anonymous):

many values of x and y can satisfy the equation!

OpenStudy (anonymous):

if your solution satisfies both(all) the given equations then and only then your solution exists.